A horizontal pipe contains water at a pressure of 130kPa flowing with a speed of 1.5m/s . The pipe narrows to one-half its original diameter.

A) What is the speed of the water? (in m/s)

B) What is the pressure of the water? (in kPa)

Thanks!
Answer #1

A) To get the speed we use the continuity equation:

A_1v_1 = A_2v_2

v_2 = \frac{A_1v_1}{A_2} = \frac{d_1^2}{d_2^2}v_1 = 4v_1 = 6.0 m/s

B) Using Bernoulli equation and considering the pipe is horizontal :
P_1 + \frac{1}{2}\rho v_1^2 = P_2 + \frac{1}{2}\rho v_2^2

solving for: P2

P_2 = P1 +\frac{1}{2}\rho (v_1^2-v_2^2)

P_2 = 147kPa